OpenStudy (anonymous):

A truck covers 40 m in 8.5 s while smoothly slowing down to a final speed of 2.8 m/s. Find the truck's original speed and it's acceleration.

OpenStudy (anonymous):

Vo = s/t = 40/ 8.5 = ....? V = Vo + at =>a = ( V - Vo ) / t